OpenStudy (anonymous):

If 0 deg < x <90 deg and 2sin^2 x + cos^2 x = 25/16. what is the value of sinx?

OpenStudy (anonymous):

2 sin^2x + cos^2x = sin^2x + sin^2x + cos^2x But we know that sin^2x + cos^2x = 1 (Pythagorean Identity) So the expression is equal to sin^2x + 1 = 25/16 Lets solve for sin x. Subtract 1 from both sides, we get sin^2x = 9/16 sin x = sqrt(9/16) sin x = 3/4So where is the sin 3/4? Use your calculator to find 49 degrees to the nearest degree.

OpenStudy (anonymous):

@Easyaspi314 Thank you!

OpenStudy (anonymous):

welcome.